### v1

# QCD corrections to the electroweak sphaleron rate

Abstract
The electroweak sphaleron rate in the high temperature phase of the Standard Model is inversely proportional to the weak-isospin conductivity. So far, only electroweak interactions were included in its computation. Here we take into account quark scattering through strong interactions at leading-log order. These reduce the quark contribution to the conductivity by up to 15%, and the total conductivity by up to 6%.
1 Introduction
The rate for electroweak baryon number violation at finite temperature $T$, also called the sphaleron rate, is a key ingredient in scenarios that could explain the baryon asymmetry of the Universe. [^3] For a recent review, see e.g. [1]. At temperatures exceeding the electroweak phase transition (or crossover) temperature $T_{c}$, the rate is larger than the Hubble rate, leading to rapid baryon number violation. Around $T_{c}$, the rate freezes out, so that baryon number is conserved from then on. At higher temperatures, the rate is determined by the non-perturbative dynamics of electroweak gauge bosons with wavelengths of order of the electroweak magnetic screening length $\ell_{\text{mag}}\sim(g_{W}^{2}T)^{-1}$, where $g_{W}$ is the weak-SU(2) gauge coupling. Due to their large occupation number $\sim T/|{\bf k}|\sim g_{W}^{-2}$, they can be treated as classical fields. Their dynamics is determined by the non-Abelian counterpart of Ampere’s law
${\bf D}\times{\bf B}=\sigma{\bf E}+\mbox{$\zeta$}\,.$ (1.1)
The right-hand side of eq. (1.1) is the weak-isospin current, which captures the medium response due to particles with hard momenta $|{\bf p}|\sim T$. The ohmic part of the current is local in the weak-electric field $\bf E$} because the mean free path of isospin-changing small-angle scattering is parametrically of order $g_{W}^{2}\log(g_{W}^{-1})T]^{-1}$. In a leading log approximation, this length scale is much smaller than the magnetic screening length. The coefficient $\sigma$ is the weak-isospin conductivity, [^4] In the context of QCD it is referred to as color conductivity. and $\zeta$ is a Gaussian white noise term generated by thermal fluctuations of the high-momentum modes.
The only length scale in this equation is $\ell_{\text{mag}}\sim(g_{W}^{2}T)^{-1}$. We can therefore estimate ${\bf D}\times{\bf B}\sim\ell_{\text{mag}}^{-2}{\bf A}$, where $\bf A$} is the weak SU(2) vector potential. The first term on the right-hand side of (1.1) acts like a damping term for the gauge fields, and determines the characteristic time scale $\tau$ of the isoelectric field ${\bf E}\sim\tau^{-1}{\bf A}$ as $\tau\sim\sigma\,\ell_{\text{mag}}^{2}$. The value of the conductivity is large, $\sigma\sim T/\log(g_{W}^{-1})$, which leads to over-damping of the gauge fields. Therefore, eq. (1.1) does not contain a “displacement current” $D_{t}{\bf E}$, which would be of order $\tau^{-2}{\bf A}$.
The hot electroweak sphaleron rate $\Gamma_{\rm sph}$, i.e. the rate of sphaleron transitions per unit time and unit volume, is parametrically of order $\Gamma_{\rm sph}\sim(\tau\ell_{\text{mag}}^{3})^{-1}$ and can be written as [1]
$\Gamma_{\rm sph}=k\frac{2\pi T}{3\sigma}\alpha_{W}^{5}T^{4}$ (1.2)
with $\alpha_{W}\equiv g_{W}^{2}/(4\pi)$. The dimensionless number $k$ has been determined through lattice simulations of eq. (1.1) [2, 3], [^5] Close to the electroweak crossover one needs to include the Higgs field in the effective infrared theory [4], for recent results, see [5]. yielding
$k=10.0\pm 0.3\,.$ (1.3)
[…]
However, the strong coupling constant $g_{S}$ is numerically larger than $g_{W}$, which may partially compensate the small-angle enhancement of the electroweak scattering. The main aim of this work is therefore to determine the quantitative impact of QCD corrections of the quark contribution to the weak-isospin conductivity $\sigma$. To incorporate the effect of QCD scattering on the electroweak sphaleron rate, it suffices to insert the new value of $\sigma$ in eq. (1.2), without altering the value of $k$.
The remainder of this work is organized as follows: In sec. 2 we review the Vlasov equations of electrodynamics describing the hard thermal loop effective theory, their non-Abelian generalization and the resulting Boltzmann equation containing the electroweak collision term. In sec. 3, we compute the QCD collision term. Sec. 4 presents the numerical inversion of the complete quark collision term to obtain the corrected isospin conductivity, and sec. 5 concludes the paper. This paper includes three appendices. Appendix A collects the coupling constants used to obtain our numerical results. A derivation of the Vlasov-Boltzmann equations from Schwinger-Dyson equations is provided by appendix. B. Appendix C details a derivation of the QCD collision term from a 2-particle-irreducible (2PI) effective action.
2 Vlasov-Boltzmann equation
To begin, we recall the Vlasov equations that govern the dynamics of an electrodynamic plasma and their relation to the physics of hard thermal loops (HTLs) in QED (cf. [7, 8]). These equations describe the coupled evolution of long wavelength ($|{\bf k}|\ll T$) thermal fluctuations of the electromagnetic field and the phase space densities of charged particles with hard ($|{\bf p}|\sim T$) momenta in a hot plasma. Such fluctuations are present even in thermal equilibrium. Due to their large occupation number, the long-wavelength fluctuations of the electromagnetic field can be treated as classical fields, while the charged particle are treated as classical particles.
The phase space densities $f_{e^{\pm}}(t,{\bf x},{\bf p})$ of positrons and electrons then evolve according the Vlasov equations [9]
$(\partial_{t}+{\bf v}\cdot\nabla)f_{e^{\pm}}\pm e({\bf E}+{\bf v}\times{\bf B})\cdot\frac{\partial f_{e^{\pm}}}{\partial{\bf p}}=0\,$, (2.1)
where $\bf v$} is the particle velocitity and $e$ is the positron charge. This equation neglects the effect of hard gauge fields, which would lead to an additional collision term.
We consider high temperatures, such that we can neglect the electron mass and the velocity is given by ${\bf v}={\bf p}/|{\bf p}|$. The electric current is
$J_{\text{em}}^{\mu}=2e\int\frac{d^{3}p}{(2\pi)^{3}}(f_{e^{+}}-f_{e^{-}})v^{\mu}\,$, (2.2)
where we have assumed that both spin polarizations give the same contribution, resulting in the factor 2. For small deviations from equilibrium $f_{e^{\pm}}=f_{\rm F}+\delta f_{e^{\pm}}$, where $f_{\rm F}({\bf p})\equiv(e^{|{\bf p}|/T}+1)^{-1}$ is the Fermi-Dirac distribution, one may linearize eq. (2.1). This gives
$(\partial_{t}+{\bf v}\cdot\nabla)\delta f_{e^{\pm}}\pm e\,{\bf E}\cdot{\bf v}\,f^{\prime}_{\rm F}$ $=0$ (2.3)
with $f^{\prime}_{\rm F}\equiv\partial f_{\rm F}/\partial|{\bf p}|$. By solving eq. (2.3) for $\delta f$, inserting the solution into the current (2.2) and Fourier transforming, one obtains
$J_{\text{em}}^{\mu}(k)=\Pi^{\mu\nu}_{\rm HTL}(k)A_{\nu}(k)\,$, (2.4)
where $\Pi^{\mu\nu}_{\rm HTL}$ is the hard thermal loop polarization tensor. This tensor is derived from thermal field theory by integrating out hard modes to obtain an effective theory for low-momentum $k^{\mu}\ll T$ modes of the electromagnetic field. In essence, the physics of classical charged particles interacting with a classical gauge field is same as the physics of hard thermal loops. The HTL effective theory is then described by eqs. (2.2) and (2.3) together with Maxwell’s equation [8]
$\partial_{\mu}F^{\mu\nu}=J_{\text{em}}^{\nu}\ .$ (2.5)
In the non-Abelian weak SU(2) theory, there is not only a HTL polarization tensor, but also HTL $n$-point functions for all $n>2$. Their generating functional is determined by non-Abelian counterparts of eqs. (2.2) and (2.3) [10], with one distribution function for each isospin-charged particle in the Standard Model. In the following, we focus on the quark contribution to the weak-isospin current
$J^{a}_{q\,\mu}=\frac{g_{W}}{2}N_{\rm fam}N_{\rm c}\int\frac{d^{3}p}{(2\pi)^{3}}\left[\delta f^{a}_{+}-\delta f^{a}_{-}\right]v_{\mu}\ ,$ (2.6)
where we have summed over colors and quark isospin doublets, resulting in the factors $N_{\rm c}=3$ and $N_{\rm fam}=3$, respectively. Note that there is no spin sum, as only left-handed quarks and right-handed antiquarks couple to the SU(2) gauge bosons. Since the doublets transform under the fundamental representation of SU(2), their distribution functions $\delta f^{a}_{\pm}$, defined as moments of quark two-point correlation functions, transform under the adjoint representation. [^6] See also the discussion in app. B for details. This behavior is similar to lepton flavor oscillations, which can be described by promoting the neutrino distribution function to a $\times3$ matrix in flavor space. The non-Abelian version of eq. (2.3) for quarks is given by [10]
$(D_{t}+{\bf v}\cdot{\bf D})\delta f_{\pm}\pm g_{W}{\bf v}\cdot{\bf E}\,f_{\rm F}^{\prime}=0\,$, (2.7)
where $D_{\mu}$ is the SU(2) covariant derivative in the adjoint represenation. The complete HTL effective theory for gauge fields with small momenta $|{\bf k}|\ll T$ is then fully described [11] by eqs. (2.6) and (2.7) along with the Yang-Mills equation
$D_{\mu}F^{\mu\nu}=J^{\nu}\ ,$ (2.8)
where the current $J^{\mu}$ is defined by adding the quark contribution (2.6) to the analogous contributions from the lepton, $W$ boson, and Higgs fields. Ultimately, we want to describe magnetic-scale gauge fields. For these, the HTL polarization tensor is of order $g_{W}^{2}T^{2}k_{0}/|{\bf k}|$, which is much larger than the tree-level terms in the propagator, unless $k_{0}\ll|{\bf k}|$. In other words, they are strongly overdamped [12], rendering time derivatives in eqs. (2.7) and (2.8) negligible.
By further intergrating out fields with momenta $k^{\mu}\sim g_{W}T$, one obtains an effective theory for the fields with momenta on the order of the magnetic screening scale $g_{W}^{2}T$. Their dynamics is governed by equations of motion that take the same form as eqs. (2.6), (2.7), and (2.8), except that there is now a gaussian white noise and a linear collision term on the right-hand side of eq. (2.7), which we denote as $\left[\dot{f}_{\pm}\right]_{\rm coll}$. The collision term is of order $g_{W}^{2}\log(g_{W}^{-1})T\delta f$, while the term ${\bf v}\cdot{\bf D}\,\delta f$ on the left-hand side of eq. (2.7) is of order $\sim g_{W}^{2}T\delta f$. In a leading-log approximation, which turns out to be valid also at next-to-leading log order, one can therefore neglect the latter term, which yields
$\pm g_{W}{\bf v}\cdot{\bf E}\,f_{\rm F}^{\prime}=\left[\dot{f}_{\pm}\right]_{\rm coll}\ .$ (2.9)
Here we have not displayed the white noise term that also results from integrating out the modes with $|{\bf k}|\gg g_{W}^{2}T$. This term gives rise to the noise term in eq. (1.1), but it is not needed for our computation of the isospin conductivity.
The linearized collision term can be viewed as a linear operator acting on $\delta f$. This operator commutes with rotations acting on the velocity variable $\bf v$}. Thus its eigenfunctions are the spherical harmonics $Y_{lm}({\bf v})$, and its eigenvalues only depend on $l$ [13]. Since the left-hand side of eq. (2.9) is a linear combination of spherical harmonics with $l=1$, only the $l=1$ sector of $\delta f_{\pm}$ is nonzero, so that
$\left[\dot{f}_{\pm}\right]_{\rm coll}$ $=-c_{W}\,\delta f_{\pm}\ ,$ (2.10)
where $c_{W}$ is minus the $l=1$ eigenvalue of the collision term. At next-to-leading log order, it is given by [6]
$c_{W}$ $=\gamma+\frac{g_{W}^{2}T}{2\pi}K\ ,$ $K$ $=3.0410\ ,$ (2.11)
where $\gamma$ is defined implicitly as a solution of the equation
$\gamma=\frac{g_{W}^{2}T}{2\pi}\log\left(\frac{m_{W}}{\gamma}\right)\ .$ (2.12)
Here
$m^{2}_{{W}}=\left(\frac{2}{3}+\frac{N_{s}}{6}+\frac{(1+N_{\rm c})N_{\rm fam}}{12}\right)g_{W}^{2}T^{2}$ (2.13)
is the $W$-boson Debye mass with $N_{s}=1$ the number of Higgs doublets in the SM. Table 1 presents a selection of values for $\gamma$. Using eq. (2.10) in eq. (2.9), one finds
$\delta f_{\pm}$ $=\frac{\mp g_{W}{\bf v}\cdot{\bf E}}{c_{W}}f_{\rm F}^{\prime}\ .$ (2.14)
Inserting this solution into eq. (2.6), one recovers the ohmic part of the current in eq. (1.1) with the isospin conductivity
$\sigma_{q}=\frac{m^{2}_{Wq}}{3\,c_{W}}\ ,$ (2.15)
where
$m^{2}_{Wq}=\frac{N_{\rm c}N_{\rm fam}}{12}g_{W}^{2}T^{2}$ (2.16)
is the quark contribution to the $W$ boson screening mass. So far, we have only accounted for quarks. Including all Standard Model fields, the conductivity turns out to be [14]
$\sigma$ $=\frac{m^{2}_{{W}}}{3\,c_{W}}\ ,$ (2.17)
which is identical to eq. (2.15), except that $m_{{\rm W}q}$ is replaced by the full screening mass $m_{W}$. This remarkable similarity arises from the simplicity of the collision term in eq. (2.10), which turns out to be the same for all Standard Model particles that carry weak isospin. In general, quarks also scatter with other isospin charged particles, so that their collision terms will “mix” quark and other distribution functions, causing them to influence each other. However, the term that mixes the distribution functions vanishes at next-to-leading log order when applied to an odd function of $\bf v$}, so this mixing does not affect the $l=1$ sector of the collision term, which is odd under ${\bf v}\to-{\bf v}$ [14]. Hence, the distribution functions effectively decouple at the level of the long-wavelength limit of the Boltzmann equation given in eq. (2.9). As a result, generalizing eq. (2.15) to include all isospin-carrying particles amounts to simply counting their isospin, color, and flavor degrees of freedom.
For our purposes, the absence of mixing terms also has the important consequence that QCD corrections to the collision term only affect the quark distribution functions, so that we do not have consider the out-of-equilibrium dynamics of the other Standard Model particles.
3 The QCD collision term
We now include the scattering of quarks off quarks and gluons via QCD interactions. These do not mix quarks with other particles that carry weak isospin. Therefore one can write the complete linearized collision term as a sum of the weak-interaction term discussed in sec. 2 and the new strong-interaction term. Like ref. [15], we work at leading logarthmic order in the strong coupling $g_{S}$. We use the collision term in the form obtained in ref. [16], and also derive it using Dyson-Schwinger equations in app. C.
(a) $qq\leftrightarrow qq$
(b) $qg\leftrightarrow qg$
(c) $qg\leftrightarrow qg$
(d) $qq\leftrightarrow gg$
Figure 1: Feynman diagrams contributing to the QCD collision term at leading-log order. The logarithmic enhancement is the result of apparent IR divergencies in the $t$-channel contribution to $2\leftrightarrow 2$ scattering processes with soft gluon or quark exchanges. Interference terms and $s$-channel contributions are IR safe and therefore only contribute at leading order.
At leading logarithmic order, only the $2\leftrightarrow 2$ processes with $t$-channel momentum exchange depicted in figure 1 contribute. If the resulting collision term were computed using tree-level propagators, the integral over the exchanged 4-momentum $q$ would contain terms that behave as $\int d|{\bf q}|/|{\bf q}|$. They are cut off by Bose-Einstein or Femi-Dirac distributions for large momenta $|{\bf q}|\gg T$, but diverge logarithmically for small $|{\bf q}|$. However, medium effects significantly modify the propagation of the exchanged quarks and gluons when $|{\bf q}|$ is of order of the Debye scale $g_{S}T$, and provide a cutoff at even smaller momenta. Due to the separation between hard ($|{\bf q}|\sim T$) and soft ($|{\bf q}|\sim g_{S}T$) scales, one then obtains contributions that are enhanced by a factor
$\int\displaylimits_{g_{S}T}^{T}\frac{d|{\bf q}|}{|{\bf q}|}=\log(g_{S}^{-1})\ .$ (3.1)
The leading-log approximations consists in keeping only these enhanced contributions. Since $s$-channel exchanges and interference terms are infrared safe, they do not contribute at leading logarithmic order.
Without loss of generality, we may consider a SU(2) electric field that it is diagonal in flavor space,
${\bf E}={\bf E}^{3}\frac{\tau^{3}}{2}\ ,$ (3.2)
where $\tau^{3}$ is the third Pauli matrix. Then the long-wavelength limit of the Boltzmann equation (2.9) for left-handed quarks of flavor $f$ takes the form
$\pm\frac{g_{W}}{2}{\bf v}\cdot{\bf E}^{3}f_{\rm F}^{\prime}=$ $-c_{W}\,\delta f_{fL}+\left[\dot{f}_{fL}\right]_{\text{strong}}\ ,$ (3.3)
where the upper and lower sign applies to up- and down-type quarks, respectively. We write the phase space densities as
$f_{\alpha}=f_{\alpha}^{\rm eq}+f_{\alpha}^{\rm eq}(1\pm f_{\alpha}^{\rm eq})\eta_{\alpha}\ ,$ (3.4)
where the index $\alpha$ denotes species and helicity. Then, one has
$\eta_{uL}=-\eta_{dL},\quad\eta_{qL}=-\eta_{\bar{q}R}\ .$ (3.5)
The right-handed quarks and left-handed antiquarks as well as the gluons remain in equilibrium, so that $\eta_{qR}=\eta_{\bar{q}L}=\eta_{g}=0$. The linearized collision term due to $t$-channel gluon exchanges, which mediate the quark-quark and quark-gluon scatterings shown in Figs. 1(a) and 1(b), can be written as [16]
[…]
where $m_{G}$ with
$m^{2}_{G}$ $=\left(\frac{N_{\rm c}}{3}+\frac{N_{\rm fam}}{3}\right)g_{S}^{2}T^{2}\$(3.7)
is the gluon Debye mass for $2N_{\rm fam}$ quark flavors. The gain terms vanish because they are they are linear operators acting on either $\eta_{g}=0$ or the sum over quark flavors $\sum_{f}\eta_{fh}$, which vanishes due to eq. (3.5) for both helicities $h=L,R$.
The analogous linearized collision term due to quark exchange, which mediate the quark-gluon scatterings shown in Fig. 1(c) and the quark-antiquark annihilation shown in Fig. 1(d), is [16]
$\left[\dot{f}_{uL}(x,{\bf p})\right]_{\text{qu}}$ $=-\frac{g_{S}^{2}\log(g_{S}^{-1})}{6\pi p_{0}}m^{2}_{{\infty}}[1+2f_{\rm B}(p_{0})]f_{\rm F}(p_{0})[1-f_{\rm F}(p_{0})]\eta_{\alpha}+\text{gain terms}\ ,$ (3.8)
where
$m^{2}_{{\infty}}$ $=\frac{N_{\rm c}^{2}-1}{8N_{\rm c}}g_{S}^{2}T^{2}$ (3.9)
is the QCD contribution to the asymptotic quark mass [17]. [^7] We have neglected the electroweak contribution to the asymtpotic mass, assuming $g_{W}\ll g_{S}$. Eq. (3.8) again contains no gain terms, which vanish as in the electroweak case due to the rotational symmetries of the collision integral.
Like the linearized electroweak collision term, the linearized QCD collision term commutes with rotations, so that $\eta_{\alpha}$ has the same angular dependence as the force term. For the left-handed up quarks we can thus write
$\eta_{uL}=\frac{g_{W}}{2T}\,{\bf v}\cdot{\bf E}^{3}\,\chi(p_{0})\ .$ (3.10)
Inserting this into eqs. (3.6) and (3.8) gives
[…]
were primes denote derivatives with respect to $p_{0}$, and
$\left[\dot{f}_{uL}(x,{\bf p})\right]_{\text{qu}}$ $=-\frac{g_{S}^{2}\log(g_{S}^{-1})}{12\pi}m^{2}_{{\infty}}\,g_{W}\,{\bf v}\cdot{\bf E}^{3}\,\frac{1-2f_{\rm B}(p_{0})}{Tp_{0}}f_{\rm F}(p_{0})[1-f_{\rm F}(p_{0})]\chi\ .$ (3.14)
Finally, insertig the sum of these two expressions into eq. (3.3), one obtains a linear differential equation
$1=c_{W}\chi-\frac{g_{S}^{2}\log(g_{S}^{-1})}{6\pi}\Bigg\{$ $m^{2}_{G}T\left[\chi^{\prime\prime}+\left(\frac{2}{p_{0}}-\frac{1}{T}[1-2f_{\rm F}(p_{0})]\right)\chi^{\prime}-\frac{2}{p_{0}^{2}}\chi\right]$
$-m^{2}_{{\infty}}\frac{1+f_{\rm B}(p_{0})}{1-f_{\rm F}(p_{0})}\frac{\chi}{p^{0}}\Bigg\}$ (3.15)
that determines the function $\chi(p_{0})$.
4 Results
To obtain the quark contribution to the weak-isospin conductivity, we have to solve eq. (3.15) numerically.
Table 1: Values for the parameter $\gamma$ entering the $l=1$ eigenvalue of the electroweak collision term $c_{W}$ in eq. (2.11) and the parameter $a$ in the differential equation (4.3). $\sigma_{\cancel{\rm QCD}}$ is the weak-isospin conductivity without QCD corrections, $\sigma_{\rm ana}$ is obtained using the analytic approximation (4.14) for $\kappa$, and $\sigma_{\rm num}$ by solving eq. (4.3) numerically.
TT/GeV $\gamma/T$ $c_{W}/T$ $a$ $\sigma_{\rm\cancel{\rm QCD}}/T$ $\sigma_{\rm ana}/T$ $\sigma_{\rm num}/T$
130 0.127 0.327 1.521 0.773 0.728 0.728
200 0.126 0.325 1.602 0.773 0.729 0.730
$10^{3}$ 0.124 0.317 1.917 0.771 0.733 0.734
$10^{8}$ 0.107 0.271 4.609 0.763 0.742 0.745
In ref. [18], it was found that the complete leading order results for various transport coefficients are well approximated by a next-to-leading log approximation that replaces
$\log(g_{S}^{-1})$ $\to\log\left(\frac{\mu_{\ast}}{m_{G}}\right)\qquad\mbox{ with }\qquad\mu_{\ast}\simeq 3T$ (4.1)
in the leading-log result. For our numerical results, we adopt the prescription (4.1). For convenience, we also define the dimensionless quantities
$a$ $=\frac{3\pi c_{W}T}{g_{s}^{2}m^{2}_{G}\log\left(\mu_{\ast}/m_{G}\right)}\ ,$ $b$ $=\frac{m^{2}_{\infty}}{m^{2}_{G}}=\frac{1}{6}\ ,$ $x$ $=\frac{p_{0}}{T}\ ,$ $F(x)$ $=\frac{c_{W}\,\chi(p_{0})}{a}\ .$ (4.2)
Hence, the differential equation (3.15) for $\chi$ turns into
$1$ $=-\frac{1}{2}F^{\prime\prime}-\left(\frac{1}{x}-\frac{1}{2}\frac{e^{x}-1}{e^{x}+1}\right)F^{\prime}+\left(a+\frac{b}{2x}\frac{e^{x}+1}{e^{x}-1}+\frac{1}{x^{2}}\right)F\ ,$ (4.3)
where the primes now denote derivatives with respect to $x$. The coefficient
$a\sim\frac{g_{W}^{2}\log(g_{W}^{-1})}{g_{S}^{4}\log(g_{S}^{-1})}$ (4.4)
parametrizes the relative size of the weak and strong collision terms. For $a\to{\infty}$, one obtains the known result for the weak-isospin conductivity that neglects strong interactions altogether. Taking $a\to 0$, one instead obtains the weak-isospin conductivity in QCD. The quark contribution to the weak isospin conductivity is now given by
$\sigma_{q}=\frac{m_{W\text{qu}}^{2}}{3\,c_{W}}\kappa\ ,$ (4.5)
where
$\kappa=\frac{6\,a}{\pi^{2}}\int\displaylimits_{0}^{\infty}dx\,\frac{x^{2}e^{x}F(x)}{(e^{x}+1)^{2}}$ (4.6)
is normalized such that one has $\kappa\to 1$ for $a\to\infty$.
4.1 Boundary conditions
To solve eq. (4.3) numerically, one has to supply boundary conditions at $x\to 0$ and $x\to\infty$. Expanding eq. (4.3) for small $x$, one obtains the approximate equation
$1$ $=-\frac{1}{2}F^{\prime\prime}-\frac{1}{x}F^{\prime}+\frac{b+1}{x^{2}}F\ ,$ (4.7)
which has the general solution
$F(x)$ $=\frac{x^{2}}{b-2}+k_{1}\,x^{n_{+}}+k_{2}\,x^{n_{-}}\ ,$ $n_{\pm}$ $=-\frac{1}{2}\pm\frac{1}{2}\sqrt{1+16(b+1)^{2}}$ (4.8)
with integration constants $k_{1,2}$. The Pauli exclusion principle implies that the quark distribution function $f_{fL}$ must remain finite as $x\to 0$, so that $k_{2}$ has to be set to zero. Since the remaining terms in eq. (4.8) vanish for small $x$, this fixes the boundary condition $F(x)\to 0$ as $x\to 0$. For large $x$, one finds the approximate equation
$1$ $=-\frac{1}{2}F^{\prime\prime}+\frac{1}{2}F^{\prime}+aF\ ,$ (4.9)
which has the general solution
$F(x)$ $=\frac{1}{a}+k_{3}\,e^{\lambda_{+}}+k_{4}\,e^{\lambda_{-}}\ ,$ $\lambda_{\pm}$ $=\frac{1}{2}\pm\frac{1}{2}\sqrt{1+16a^{2}}$ (4.10)
with integration constants $k_{3,4}$. Like before, the Pauli exclusion principle implies that $k_{3}$ has to be set to zero to ensure that $f_{fL}$ remains finite. Therefore, the second boundary condition is $F(x)\to 1/a$ for $x\to\infty$.
To obtain the numerical solution for the conductivity shown in fig. 3, we have integrated (4.3) from $x_{\text{min}}=10^{-3}$ to $x_{\text{max}}=30$ while enforcing boundary conditions $F(x_{\text{min}})=0$ and $F(x_{\text{max}})=1/a$. We checked that varying $x_{\text{min}}$ in the range from $10^{-4}$ to $10^{-2}$ and $x_{\text{max}}$ from 2020 to 4040 barely affects the final result.
4.2 Variational approach and analytic approximation
Solving the differential equation for $F$ is equivalent to solving a variational problem $\delta Q=0$ with a functional $Q[F]$ containing terms that are either linear and quadratic in $F$ [19].
Figure 2: Solutions for $F(x)$ at $T/M_{Z}=10$. The solid line shows the results from direct numercal integration of eq. (4.3), which is indistinguishable from the variational result with $N=10$ basis functions. The dash-dotted line shows the $1/a$ expansion in eqs. (4.16), and the dashed line the result of the $N=2$ variational ansatz.
This is possible because the linearized collision term can be written as a functional derivative of a quadratic form [16]. In our case, we have to maximize the integral
$Q[F]=\int\limits_{0}^{\infty}\frac{x^{2}e^{x}}{(e^{x}+1)^{2}}\left\{F-\frac{1}{4}F^{\prime}{}^{2}-\left(\frac{a}{2}+\frac{b}{4x}\frac{e^{x}+1}{e^{x}-1}+\frac{1}{2x^{2}}\right)F^{2}\right\}\ .$ (4.11)
We use the Ansatz
$F(x)=\sum_{m=1}^{N}c_{m}\varphi_{m}(x)\ ,$ (4.12)
with $N$ basis functions [15]
$\varphi_{m}(x)=\frac{x^{m}}{(1+x)^{N-1}}\qquad(m=1,\ldots,N)\ .$ (4.13)
The maximum of $Q$ is then found by solving $\partial Q/\partial c_{m}=0$, which determines the $c_{m}$. Already with $N=2$ basis functions, we find that the result for $F(x)$ is a good approximation of the direct numerical solution of eq. (4.3) for not too large values of $x$. For example, when $a=1.8$ and $1<x<5$, it differs from the direct numerical solution by less than 1.5%. Using the value of $c_{1}$ that follows from the variational approach and setting $c_{2}\to 0$, so that only the function $\varphi_{1}(x)=x/(1+x)$ from this basis contributes to $\kappa$, we find the analytic approximation
$\kappa=\alpha\frac{a}{a+\beta}\qquad\mbox{ with }\qquad\alpha=0.9767,\quad\beta=0.2107\ .$ (4.14)
In the temperature range shown in Fig. 3, this expression deviates from the numerical result by less than 1.1%.
4.3 Weak-isospin conductivity
Figure 3: Quark contribution to the isospin conductivity with and without the strong collision term. The ratio $\kappa$ is defined in eq. (4.6). Solid lines show results from direct numercial integration of eq. (4.3), which are identical to the variational results with $N=10$ basis functions. Dash-dotted lines show results of the $1/a$ expansion given in eqs. (4.16) and (4.17), and dotted lines show the analytic approximation eq. (4.14). The gauge couplings $g_{S}$ and $g_{W}$ are evaluated at the renormalization scale $\mu=\pi T$, see app. A for details.
Figure 3 depicts the quark contribution to the weak-isospin conductivity for temperatures above the electroweak crossover. Overall, QCD interactions decrease the quark contribution to the conductivity by about 10 to 15% and are most relevant for low temperatures. The full lines are obtained by direct numerical integration of eq. (4.3). This treats the electroweak and the QCD collision terms on equal footing, and corresponds to the power counting
$g_{W}^{2}\log(g_{W}^{-1})\sim g_{S}^{4}\log(g_{S}^{-1})\ ,$ (4.15)
or equivalently $a\sim 1$. Alternatively, one may want to count $g_{S}^{4}\ll g_{W}^{2}$ or $a\gg 1$, which is valid when the QCD collision term is a small correction to the electroweak one. Then, we can expand
$F(x)$ $=\frac{1}{a}-\frac{1}{a^{2}}\left(\frac{1}{x^{2}}+\frac{b}{2x}\frac{e^{x}+1}{e^{x}-1}\right)+\mathrm{O}(a^{-3})\ .$ (4.16)
The order $a^{-2}$ term gives a finite correction to the integral in eq. (4.6), [^8] Even higher orders in $a^{-1}$ would give rise to divergent integrals. This is, however, no reason for concern, because there are other higher-order contributions that we have not accounted for. so that one obtains
$\kappa$ $=1-\left(\frac{1}{\pi^{2}}+\frac{b}{8}\right)\frac{3}{a}+\mathrm{O}(a^{-2})\ .$ (4.17)
Treating the QCD collision term as a perturbation in this way, one obtains the dash-dotted lines. The result deviates from the full result by another 5 to 10 %. Thus, it appears more sensible to fully include the QCD interactions. For this, eq. (4.14) (drawn in dotted lines) provides a simple analytic approximation that works astonishingly well, deviating form the numerical result by less than 1.1%.
5 Summary and conclusions
In the high-temperature phase of the Standard Model, when the expectation value of the Higgs field vanishes, the dynamics of the weak SU(2) gauge fields is non-perturbative. It determines the so-called hot sphaleron rate, which in turn governs the rate for baryon number violation. The gauge field dynamics is overdamped, and the damping coefficient is the weak isospin conductivity. We have computed QCD corrections to this conductivity at leading lograthmic order, and find that they reduce the value of the conductivity by 6% near the electroweak crossover. As shown in table 1, the size of the corrections decreases at higher temperatures. The corrected value can directly be used to update of the hot sphaleron rate. Including QCD corrections, the complete isospin conductivity is given by
$\sigma=\frac{1}{3c_{W}}\Big[m^{2}_{{W}}-(1-\kappa)m^{2}_{Wq}\Big]\ ,$ (5.1)
where $\kappa$ parametrizes the impact of strong interactions. A simple yet accurate analytic approximation for $\kappa$ is given by eq. (4.14).
Acknowledgments
We thank G.D. Moore for insightful comments. DB and PK acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 ’Strong-interaction matter under extreme conditions’– project number 315477589 – TRR 211. PK was also supported by the Swiss National Science Foundation (SNSF) under grant P500PT-217885.
Appendix A Coupling constants
To numerically evaluate the quark contribution to the isospin conductivity, one has to specify values for the coupling constants $g_{S}$ and $g_{W}$, which entails choosing a renormalization scale $\mu$. At one loop, the couplings are [^9] See, e.g., [20].
$\frac{1}{g_{W}^{2}}$ $=\frac{1}{g_{W}^{2}(m_{Z})}+\frac{\beta_{W}}{8\pi^{2}}\log\left(\frac{\mu}{m_{Z}}\right)\ ,$ (A.1)
$\frac{1}{g_{S}^{2}}$ $=\frac{1}{\big[g_{S}^{(5)}(m_{Z})\big]^{2}}+\frac{1}{12\pi^{2}}\log\left(\frac{m_{t}}{m_{Z}}\right)+\frac{\beta_{S}}{8\pi^{2}}\log\left(\frac{\mu}{m_{Z}}\right)\,$ (A.2)
where $m_{Z}=91.1880\pm 0.0020\,\mathrm{GeV}$ is the $Z$ Boson mass and $g_{W}(m_{Z})=0.65163\pm 0.00008$ the weak gauge coupling at $\mu=m_{Z}$ in the modified minimal subtraction ($\overline{\text{MS}}$) scheme [21]. Furthermore, $g_{S}^{(5)}(m_{Z})=1.218\pm 0.005$ the strong gauge coupling at zero temperature and at $\mu=m_{Z}$, i.e., below the top quark mass $m_{t}=162.69\pm 0.006\,\mathrm{GeV}$, where only 5 quark flavors are active. [^10] Ref. [21] reports a $1.7\sigma$ discrepancy between the pole mass from direct measurements at the LHC $m_{t}^{\text{pole}}=172.52\pm 0.44\,\mathrm{GeV}$ and the mass from electroweak precision data $m_{t}^{\text{pole}}=176.12\pm 1.9\,\mathrm{GeV}$. To compute $g_{S}$, we use the $\overline{\text{MS}}$ mass $m_{t}=162.69\pm 0.006\,\mathrm{GeV}$ that one obtains by converting the direct measurement of the pole mass [21]. We checked that using the pole-mass from electroweak precision data instead does not affect the first three significant digits of $\sigma/T$ shown in tab. 1. Considering temperatures above the electroweak crossover, we use the value of the strong coupling with $2N_{\rm fam}=6$ active quark flavors, which differs from the value with 5 flavors by the matching term proportional to $\log(m_{t}/m_{Z})$. The beta function coefficients are
$\beta_{S}$ $=\frac{11}{3}N_{\rm c}-\frac{4}{3}N_{\rm fam}\ ,$ $\beta_{W}$ $=\frac{22}{3}-\frac{1}{6}-\frac{4}{3}N_{\rm fam}\ .$ (A.3)
Since the quark contribution to the weak-isospin current 2.6 is dominated by hard momenta $p\sim\pi T$, we choose a renormalization scale $\mu=\pi T$. This results in temperature dependent values of the coupling constants $g_{S}$ and $g_{W}$, which is in fact the dominant source of the temperature dependence for the isospin conductivity shown in fig. 3.
Appendix B Vlasov-Boltzmann from Schwinger-Dyson equations
In this appendix, we use a 2PI effective approach to derive Vlasov-Boltzmann equations for the left-handed quark distribution functions $f_{fL}$ in the presence of a soft and classical non-Abelian gauge field $A_{\mu}$ from first principles of non-equilibrium quantum field theory. The derivation of Boltzmann equations via the 2PI approach is well understood (see e.g. [22, 23, 24, 25, 26, 27]), but it is comparatively less established how to modify it to account for the classical gauge field while ensuring gauge covariance. That being said, gauge covariant transport equations for non-Abelian gauge theories were previously derived in [28] and [11], and used in [11] to derive Vlasov equations that are valid at leading order in a hard thermal loop expansion.
We follow the same overall approach as in [28], and consider the time-evolution of the two point functions
$i\hat{S}^{ab}(u,v)$ $=W(x,u)\,\langle\mathscr{T}_{C}\big\{q^{a}_{fL}(u)\otimes\overline{q}^{b}_{fL}(v)\big\}\rangle\,W(v,x)\ ,$ $x$ $=\frac{u+v}{2}\ ,$ (B.1)
where the $q_{fL}$ are left-handed quark fields of flavor $f$, $\otimes$ is a tensor product in Dirac, isospin, and color space, and $\langle\cdot\rangle$ a quantum-statistical average. The time path-ordering operator $\mathscr{T}_{C}$ is defined such that fields are evaluated in sequence along the closed time-path $C$, which begins at some initial time $t_{i}$ in the far past, $t_{i}\to$-$\infty$, and goes forward until reaches some final time $t_{f}$ in the far future, $t_{f}\to+\infty$, where it turns around and returns back to $t_{i}$. Therefore, the closed time path has two separate branches, and the indices $a,b=\pm 1$ denote that the corresponding quark field is to be evaluated along either the forward (“++”) or backward (“-”) branch. If both fields are evaluated along the same branch, $\mathscr{T}_{C}$ reduces to the standard (anti-)time ordering operators familiar form zero-temperature QFT. Finally, the Wilson lines
$W\big(x+\frac{r}{2},x\big)$ $=\mathscr{P}\exp\left(ig_{W}\int\displaylimits_{0}^{1}\text{d}t\,\frac{r^{\mu}}{2}A_{\mu}\big(x+\frac{r}{2}t\big)\right)\ ,$ (B.2)
where $A_{\mu}$ is a soft and classical gauge field and $\mathscr{P}$ is a path-ordering operator, ensure that $i\hat{S}(u,v)$ transforms under the adjoint representation at $x$. Here and in the following, we suppress isospin and color indices for the sake of readability, but it should be remembered that $A_{\mu}$, $W$, and $i\hat{S}^{ab}$ are matrices in isospin and color space. On the other hand, there is no flavor mixing because we neglect Yukawa interactions and work in the symmetric phase of the Standard Model.
We primarily focus on the Wightman functions $\hat{S}^{>}=\hat{S}^{-+}$ and $\hat{S}^{<}=\hat{S}^{+-}$, which encode information about the quark distribution functions
$f_{fL}({\bf p})$ $\equiv-\int\frac{\text{d}p_{0}}{2\pi}\Theta(p_{0})\ \text{tr}_{D}\left[\gamma^{0}iS^{<}\right]\ ,$ $f_{\overline{f}R}(-{\bf p})$ $\equiv\int\frac{\text{d}p_{0}}{2\pi}\Theta(-p_{0})\ \text{tr}_{D}\left[\gamma^{0}iS^{>}\right]\ ,$ (B.3)
where $f_{\overline{f}R}$ is the distribution function of a right-handed anti-quark of flavor $f$, and
$iS^{ab}(x,p)$ $=\int\text{d}^{4}r\,e^{inr}\,i\hat{S}^{ab}\big(x+\frac{r}{2},x-\frac{r}{2}\big)$ (B.4)
are the Wightman functions in Wigner-space. Since the quarks are massless, we may parametrize the propagators as
$iS^{ab}$ $=P_{L}\gamma^{\mu}P_{R}iS^{ab}_{\mu}\ .$ (B.5)
We also need the spectral, advanced and retarded combinations
$S^{A}$ $=\frac{1}{2i}\left(S^{a}-S^{r}\right)\ ,$ $S^{a,r}$ $=S^{++}-S^{>,<}=S^{<,>}-S^{--}\ .$ (B.6)
To start, let us neglect the classical field $A_{\mu}$, so that the Wilson lines are trivial. While it is well known how to derive Boltzmann equations in this limit, we provide a short outine of the approach to better set up the subsequent inclusion of the background field $A_{\mu}$. At leading order in a gradient expansion and to any loop order, it can be shown that the Wightman functions obey the kinetic equation [22, 23, 26]
$i\partial^{\mu}\,iS_{\mu}^{>,<}=\Sigma_{\mu}^{r}\,iS^{>,<\,\mu}+i\Sigma_{\mu}^{>,<}\,S^{a\,\mu}-iS_{\mu}^{>,<}\,\Sigma^{a\,\mu}-S_{\mu}^{r}\,i\Sigma^{>,<\,\mu}\ ,$ (B.7)
where the self-energies in Wigner space
$i\Sigma_{\mu}^{ab}(x,p)$ $\equiv\frac{1}{2}\int\text{d}^{4}r\,e^{ipr}\,\text{tr}_{D}\left[P_{L}\gamma_{\mu}i\hat{\Sigma}^{ab}\big(x+\frac{r}{2},x-\frac{r}{2}\big)\right]$ (B.8)
and position space
$i\hat{\Sigma}^{ab}(u,v)$ $\equiv ab\frac{\delta\,i\mathrm{\Gamma}_{2}}{\delta\,iS^{ba}(v,u)}$ (B.9)
are consistently defined as variations of the interaction part of the 2PI effective action $\mathrm{\Gamma}_{2}$ [29, 23, 25]. The advanced, retarded, and spectral selfenergies are defined analogously to eq. (B.6). At linear order in deviations from equilibrium, the projected self-energies $i\Sigma^{ab}_{\mu}$ and propagators $iS^{ab}_{\mu}$ commute, yielding
$\frac{1}{2}\partial^{\mu}\,iS_{\mu}^{>,<}=-\Sigma_{\mu}^{\mathrm{A}}\,iS^{>,<\,\mu}+i\Sigma_{\mu}^{>,<}\,S^{\mathrm{A}\,\mu}\ .$ (B.10)
To make contact with kinetic theory, we evaluate this equation at leading order in a perturbative loop expansion, inserting the tree-level Kadanoff-Baym Ansatz
$[iS_{\mu}^{>}]_{\text{tree}}$ $=p_{\mu}\,2\pi\,\text{sign}(p_{0})\delta(p^{2})\left[(1-f_{fL}({\bf p}))\Theta(p_{0})+f_{\overline{f}R}(-{\bf p})\Theta(-p_{0})\right]\ ,$ (B.11a)
$[iS_{\mu}^{<}]_{\text{tree}}$ $=p_{\mu}\,2\pi\,\text{sign}(p_{0})\delta(p^{2})\left[-f_{fL}({\bf p})\Theta(p_{0})-(1-f_{\overline{f}R}(-{\bf p}))\Theta(-p_{0})\right]\ .$ (B.11b)
Hence, multiplying both sides of eq. (B.10) by $\Theta(p_{0})$, integrating over $p_{0}$, adding up the >> and << cases, and linearizing in the deviation from equilibrium, one obtains
$v^{\mu}\,\partial_{\mu}f_{fL}({\bf p})$ $=\left[\dot{f}_{fL}\right]_{\text{coll}}\ ,$ (B.12)
where $v^{\mu}\equiv(1,\nicefrac{{p^{i}}}{{p_{0}}})$ and the collision term is given as
$\left[\dot{f}_{fL}\right]_{\text{coll}}$ $=v^{\mu}\left[-2f_{fL}({\bf p})\Sigma_{\mu\,\text{eq}}^{\mathrm{A}}+\delta\Sigma_{\mu}^{\mathrm{A}}(1-2f_{\rm F}(p_{0}))-i\delta\Sigma_{\mu}^{+}\right]\ .$ (B.13)
It remains to include the classical field $A_{\mu}$. Since the leading collision term on the right-hand side of eq. (B.12) is already linear in the deviation from equilibrium, we only need to modify the Liouville operator $v^{\mu}\partial_{\mu}f_{fL}$ on the left-hand side. This can be done by considering the time-evolution of the Wightman functions in a free theory, where they obey the Dirac equation
[…]
Keeping in mind that $x=\nicefrac{{(u+v)}}{{2}}$, the derivatives of the Wilson lines are [28]
$i\partial_{\mu}^{u}\,W(u,x)$ $=-gA_{\mu}(u)W(u,x)+\frac{g}{2}W(u,x)A_{\mu}(x)-G_{1}(u,x)\ ,$ (B.15a)
$i\partial_{\mu}^{u}\,W(x,v)$ $=-\frac{g}{2}A_{\mu}(x)W(x,v)-G_{0}(x,v)\ ,$ (B.15b)
where
$G_{n}(u,x)$ $=g\int\displaylimits_{0}^{1}\text{d}t\,\frac{n+t}{2}W(u,u_{t})\,\dot{u}_{t}^{\alpha}F_{\mu\alpha}(u_{t})\,W(u_{t},x)\ ,$ $u_{t}$ $=x+(u-x)t\ .$ (B.16)
Hence, defining r=uvr=u-v, one obtains
[…]
where
$iD_{\mu\,\text{adj}}(x)\,i\hat{S}^{>,<}$ $\equiv i\partial^{x}_{\mu}\,i\hat{S}^{>,<}+g\big[A_{\mu}(x),i\hat{S}^{>,<}\big]$ (B.18)
is now the covariant derivative in the adjoint representation. Since the gauge field is soft, derivatives acting on the field strength tensor $F_{\mu\alpha}(u_{t})$ scale as $D_{t}F_{\mu\alpha}\sim g^{4}\ln(g^{-1})F_{\mu\alpha}$ and $D_{i}F_{\mu\alpha}\sim g^{2}F_{\mu\alpha}$. Working at leading order in the coupling, we can therefore neglect these gradients and replace $F_{\mu\alpha}(u_{t})\to F_{\mu\alpha}(x)$ in $G_{n}$. For the same reason, we may also neglect the gauge fields in the remaining Wilson lines, which gives
[…]
To proceed, we transform to Wigner space, consider the imaginary part of the resulting equation, and take a Dirac trace, which yields the kinetic equation
$\frac{1}{2}iD_{\text{adj}}^{\mu}(x)\,iS_{\mu}^{>,<}(x,p)+i\frac{g}{4}\,\big\{F^{\mu\alpha}(x),\partial^{p}_{\alpha}\,iS^{>,<}_{\mu}(x,p)\big\}$ $=0\ .$ (B.20)
This is the equivalent of eq. (B.10) for a free quark in the presence of a soft background field. Proceeding as we have before for eq. (B.7) (i.e. inserting the tree-level expressions (B.11) into eq. (B.20), multiplying by $\Theta(p_{0})$, integrating over $p_{0}$, adding up the >> and << cases, and finally linearizing in the deviation from equilibrium), one recovers the Vlasov equation
$v_{\mu}D_{\text{adj}}^{\mu}(x)\,f_{fL}({\bf p})+gE_{i}(x)v^{i}f_{\rm F}^{\prime}(p_{0})$ $=0\ ,$ $E_{i}$ $=F_{0i}\ .$ (B.21)
Comparing this result with the Boltzmann equation (B.12), we see that the soft classical field can be included via the replacement
$v^{\mu}\partial_{\mu}\,f_{fL}({\bf p})$ $\to v_{\mu}D_{\text{adj}}^{\mu}(x)\,f_{fL}({\bf p})+gE_{i}(x)v^{i}f_{\rm F}^{\prime}(p_{0})\ .$ (B.22)
Indeed, keeping track of both the classical field and the interacting part of the 2PI effective action, which generates the collision term, and repeating the same derivation, one obtains the linearized non-Abelian Vlasov-Boltzmann equation
$v_{\mu}D_{\text{adj}}^{\mu}(x)\,f_{fL}({\bf p})+gE_{i}(x)v^{i}f_{\rm F}^{\prime}(p_{0})$ $=\left[\dot{f}_{fL}\right]_{\text{coll}}\ .$ (B.23)
Following the discussion in sec. 2, we further neglect the $v_{\mu}D^{\mu}f_{fL}$ term when computing QCD corrections to the isospin conductivity at leading-logarithmic accuracy. Hence, inserting the expression for $E_{i}$ in eq. (3.2) into eq. (B.23), one recovers eq. (3.3).
Appendix C QCD collision term from 2PI effective action
In this appendix, we compute the QCD collision term defined in eq. (B.13).
$\supset$
Figure 4: Leading logarithmic contributions to the quark selfenergy. The cut lines represent Wightman functions while uncut lines represent (anti-)time ordered propagators. The thick double lines on the left-hand side denote exact propagators, which can be expanded perturbatively to obtain the two diagrams on the right-hand side, thus yielding the relevant logarithmically enhanced contributions.
To do so, one has to supply an apropriate expression for the self-energy (B.8). At leading logarithmic order, it is sufficient to consider the resummed one-loop diagram shown in Fig. 4(a),
[…]
$=-\frac{4g_{S}^{2}}{3}\int\frac{\text{d}^{4}k}{(2\pi)^{4}}\frac{\left(2p^{\nu}g^{\mu\rho}-p^{\mu}g^{\nu\rho}\right)}{p_{0}}\,iS^{>,<}_{\mu}(k)\,i\Delta^{>,<}_{\nu\rho}(p-k)\ ,$ (C.1)
where $iS^{>,<}(k)$ and $i\Delta^{>,<}_{\nu\rho}$ are exact quark and gluon Wightman functions. These functions can be expanded perturbatively to obtain the two unresummed two-loop diagrams shown in Figs. 4(b) and 4(c). Using the optical theorem, it can be shown that these diagrams encode the relevant logarithmically enhanced t-channel contributions from $2\leftrightarrow 2$ scattering processes shown in Fig. 1.
To extract these contributions, one has to insert appropriate tree-level and one-loop expressions for the Wightman functions into eq. (C.1) while neglecting IR safe terms that do not contribute to the leading-log result. Since gluons carry no isospin charge, their propagator does not depart from thermal equilibrium. This greatly simplifies the computation, as one can use well-known thermal equilibrium expressions.
Gluon exchanges
First, we consider the gluon exchange contribution depicted in Fig. 4(b). This diagram encodes the same physics as the $qq\leftrightarrow qq$ and $qg\leftrightarrow qg$ diagrams depicted in Figs. 1(a) and 1(b). To extract this contribution, we replace
$iS_{\mu}^{>,<}$ $\to[iS_{\mu}^{>,<}]_{\text{tree}}\ ,$ $i\Delta^{>,<}_{G\,\nu\rho}$ $\to[i\Delta^{>,<}_{G\,\nu\rho}]_{\text{one-loop}}\ ,$ (C.2)
where the tree-level quark propagator is the same as in eq. (B.11). The thermal equilibrium gluon propagator
$i\Delta^{<}_{G\,\nu\rho}(q)$ $=2f_{\rm B}(q_{0})\,\rho_{\mu\nu}(q)\ ,$ $i\Delta^{>}_{G\,\nu\rho}(q)$ $=2(1+f_{\rm B}(q_{0}))\,\rho_{\mu\nu}(q)\ ,$ (C.3)
is fully determined by the spectral function $\rho_{\mu\nu}$. In a general covariant gauge, it is given as [30, 31]
$[\rho_{\mu\nu}(q)]_{\text{one-loop}}$ $=\frac{-P^{T}_{\mu\nu}\mathrm{\Pi}^{T}-P^{L}_{\mu\nu}\mathrm{\Pi}^{L}}{q^{4}}-\alpha\frac{q_{\mu}q_{\nu}}{q^{2}}\pi\,\text{sign}(q_{0})\,\delta(q^{2})\ ,$ (C.4)
where $\alpha$ is a gauge-fixing parameter, $\mathrm{\Pi}^{L/T}$ are transverse and longitudinal spectral self-energies, while
$P^{T}_{\nu\rho}$ $=\delta_{\nu}^{i}\delta_{\rho}^{j}\left(g_{ij}+\frac{g_{00}q_{i}q_{j}}{{\bf q}^{2}}\right)\ ,$ $P^{L}_{\nu\rho}$ $=-\frac{(q^{2}g_{0\nu}-q_{0}q_{\nu})(q^{2}g_{0\rho}-q_{0}q_{\rho})}{q^{2}{\bf q}^{2}}$ (C.5)
are the corresponding projection operators. Focussing on t-channel ($q^{2}<0$) contributions and dropping terms that are IR safe in the $|{\bf q}|\to 0$ limit, we replace the full self-energies by their HTL expressions [30]
$\mathrm{\Pi}^{L}$ $\overset{\text{HTL}}{=}-2\mathrm{\Pi}^{T}\overset{\text{HTL}}{=}\frac{\pi}{2}m^{2}_{G}\frac{q_{0}}{|{\bf q}|}\left(1-\frac{q_{0}^{2}}{{\bf q}^{2}}\right)\ .$ (C.6)
Inserting the one-loop expressions (C.3) and (C.4) for the gluon propagator and the tree-level quark propagator (B.11) into the eq. (C.1), one obtains a collision term that sums all contributions from $2\leftrightarrow 2$ and kinematically allowed $1\leftrightarrow 3$ processes mediated by a gluon exchange while neglecting interference terms. To isolate only t-channel exchanges, we require that the energy $k_{0}$ of the quark propagator in eq. (C.1) is positive. Hence, using the Ansatz (3.4) to parametrize the deviation from equilibrium, and dropping IR safe contributions, one obtains the collision term
$\left[\dot{f}_{fL}({\bf p})\right]_{\text{gl}}$ $=\frac{m^{2}_{G}g_{S}^{2}}{4\pi^{2}}\int\frac{\text{d}^{3}{\bf q}}{|{\bf q}|^{5}}(1-f_{\rm F}(p_{0}))f_{\rm F}(p_{0}-q_{0})q_{0}f_{\rm B}(q_{0})\left[\eta_{fL}({\bf p}-{\bf q})-\eta_{fL}({\bf p})\right]$
$=\frac{m^{2}_{G}g_{S}^{2}T}{6\pi}\delta^{ij}\partial_{p_{i}}\left[(1-f_{\rm F}(p_{0}))f_{\rm F}(p_{0})\partial_{p^{j}}\,\eta_{fL}({\bf p})\right]\int\frac{\text{d}|{\bf q}|}{|{\bf q}|}+\text{ IR safe}\ ,$ (C.7)
where
$q_{0}$ $\equiv|{\bf p}|-|{\bf p}-{\bf q}|=(\hat{p}\hat{q})\frac{|{\bf q}|}{p_{0}}+O(\nicefrac{{|{\bf q}|^{2}}}{{p_{0}^{2}}})$ (C.8)
is the energy of the exchanged gluon. Note that this expression is manifestly independent of the gauge fixing parameter $\alpha$. To finally extract the leading-logarithmic contribution, we follow e.g. [14, 15] and impose the IR and UV cutoff
$\int\frac{\text{d}|{\bf q}|}{|{\bf q}|}$ $\to\int\displaylimits_{g_{S}T}^{T}\frac{\text{d}|{\bf q}|}{|{\bf q}|}=\log(g_{S}^{-1})\ .$ (C.9)
Hence, inserting this replacement into eq. (C.7), we recover eq. (3.6) for the gluon exchange contribution to the QCD collision term, which is also given e.g. in [16].
Quark exchanges
Next, we consider the quark exchange contribution depicted in Fig. 4(c). Up to interference terms, which are captured by separate two-loop diagrams, this diagram encodes the same physics as the $qg\leftrightarrow qg$ and $qq\leftrightarrow gg$ diagrams depicted in Figs. 1(c) and 1(d). To extract this contribution, we replace
$iS_{\mu}^{>,<}$ $\to[iS_{\mu}^{>,<}]_{\text{one-loop}}\ ,$ $i\Delta^{>,<}_{\nu\rho}$ $\to[i\Delta^{>,<}_{\nu\rho}]_{\text{tree}}\ ,$ (C.10)
where the tree-level gluon propagator is the same as in eq. (C.3), except with the one loop spectral function replaced by the well-known tree-level expression
$[\rho_{\mu\nu}(q)]_{\text{tree}}$ $=\left(-g_{\mu\nu}+(1-\alpha)\frac{q_{\mu}q_{\nu}}{q^{2}}\right)\pi\,\text{sign}(q_{0})\,\delta(q^{2})\ .$ (C.11)
The one-loop quark propagator
$[iS_{\mu}^{>,<}(q)]_{\text{one-loop}}$ $=\frac{2q^{\mu}q^{\sigma}-q^{2}g^{\mu\sigma}}{q^{4}}[i\Sigma_{\sigma}^{>,<}(q)]_{\text{one-loop}}$ (C.12)
is defined recursively in terms of the strict one-loop contribution to the quark self-energy (C.1), which is given by replacing both the quark propagator and gluon spectral function by their respective tree-level expressions eqs. (B.11) and (C.11),
$i\Sigma_{\sigma}^{>,<}(q)]_{\text{one-loop}}=-\frac{4g_{S}^{2}}{3}\int\frac{\text{d}^{4}k}{(2\pi)^{4}}\left(2g^{\sigma\alpha}g^{\beta\gamma}-g^{\sigma\gamma}g^{\alpha\beta}\right)\\ \times[iS^{>,<}_{\gamma}(k)]_{\text{tree}}\,[i\Delta^{>,<}_{\alpha\beta}(q-k)]_{\text{tree}}\ .$ (C.13)
Inserting the expression in eqs. (C.11) and (C.12) into (C.1), and collecting all terms with a non-trivial Lorentz structure, one has to evaluate the object
$\frac{1}{q^{4}}(2q^{\mu}q^{\sigma}-q^{2}g^{\mu\sigma})Q_{\mu}(p,q)Q_{\sigma}(k,q)\ ,$ (C.14)
where $p$ and $k$ are the momenta of the two quarks in the initial or final state, $q$ the exchanged momentum, (pq)(p-q) and (kq)(k-q) the resulting gluon momenta. The function
$Q^{\mu}(p,q)$ $\equiv(2g^{\mu\rho}p^{\nu}-p^{\mu}g^{\nu\rho})\left(-g_{\nu\rho}+(1-\alpha)\frac{(p-q)_{\nu}(p-q)_{\rho}}{(p-q)^{2}}\right)$ (C.15)
depends on the gauge-fixing parameter $\alpha$. This is to be expected because we have neglected various two-loop diagrams that capture interference terms, and that would have to be included for a complete leading order computation of the collision term. Therefore, the object given in (C.14) can only be gauge parameter invariant at leading-logarithmic order, i.e. after neglecting IR safe contributions and fixing the signs of the quark and gluon momenta such that $k_{0}(k_{0}-q_{0})>0$ and $(p_{0}-q_{0})>0$. This gives
$\frac{1}{q^{4}}(2q^{\mu}q^{\sigma}-q^{2}g^{\mu\sigma})Q^{\mu}(p,q)Q^{\sigma}(k,q)$ $\to-\frac{(pk)}{q^{2}}+\text{ IR safe}\ .$ (C.16)
In this limit, inserting the Ansatz (3.4), and using the cutoff prescription (C.9), one recovers the collision term
$\left[\dot{f}_{fL}(x,{\bf p})\right]_{\text{qu}}$ $=-\frac{g_{s}^{2}m_{\infty}^{2}\ln(g_{s}^{-1})}{6\pi p_{0}}(1+2n_{B}(p_{0}))n_{F}(p_{0})(1-n_{F}(p_{0}))\eta_{fL}({\bf p})\ ,$ (C.17)
where a gain term that also depends on $\eta_{fL}({\bf p})$ vanishes since the departure from equilibrium is odd under ${\bf p}\to-{\bf p}$. This is the expression we use in eq. (3.8), and it also is given in [16].

### v2

@@
 is the gluon Debye mass for $2N_{\rm fam}$ quark flavors.
-The gain terms vanish because they are they are linear operators acting on either $\eta_{g}=0$ or the sum over quark flavors $\sum_{f}\eta_{fh}$, which vanishes due to eq.
+The gain terms vanish because they are linear operators acting on either $\eta_{g}=0$ or the sum over quark flavors $\sum_{f}\eta_{fh}$, which vanishes due to eq.
 (3.5) for both helicities $h=L,R$.
@@
 where $\kappa$ parametrizes the impact of strong interactions. A simple yet accurate analytic approximation for $\kappa$ is given by eq. (4.14).
+Our results show that QCD interactions give rise to a relatively small correction to the weak-isospin conductivity, but also that the correction is big enough that one should treat the QCD interactions on the same level as the weak interactions when solving the Boltzmann equation.
+We have worked at leading-log order both in the weak and in the strong coupling.
+For the collision term this involved incorporating only the contributions of order $g_{W}^{2}\log(g_{W}^{-1})$ and $g_{S}^{4}\log(g_{S}^{-1})$.
+Going beyond this approximation should be relatively straightforward for the QCD contribution.
+A first step would be to include the non-logarithmic leading order contributions from $2\leftrightarrow 2$ scatterings and near-collinear in-medium emissions, which have been computed for QCD transport coeffcients [18].
+Additionally, one could also incorporate order $g_{S}$ corrections [20], which were computed for QCD transport coefficients in Ref.
+[16].
+It is less clear how to go beyond the leading-log approximation in the purely electroweak sector.
+While it would be possible to include the relevant corrections to the collision term, [^9] For example, one could incorporate order $g_{S}^{2}$ thermal quark masses in the weak collision term, which would result in $g_{W}^{2}g_{S}^{2}$ contributions.
+However, there are additional scattering process that contribute at the same order, and that would need to be included on the same footing to achieve a fully consistent result.
+It is also important to note that these contributions are of higher order than the order $g_{W}^{2}$ part of the collision term without logarithmic enhancement that we have neglected.
+one could then no longer neglect the Liouville term on the left-hand side of the Boltzmann equation (2.9) [21, 22], which significantly complicates the computation.
+This would also generate additional operators on the right-hand side of Eq.
+(1.1), so that the impact of higher momentum modes on the evolution of the magnetic-scale gauge fields is no longer characterized by a single conductivity.
 Acknowledgments